OpenStudy (anonymous):

write the trig expression in terms of Sine and cosine, then simplify : tan^2(x)-Sec^2(x)

OpenStudy (anonymous):

hold on:

OpenStudy (anonymous):

where do I go from there?

OpenStudy (anonymous):

tan = sinx/cosx, so: tan^2 = (sinx/cosx)(sinx/cosx) = sin^2x/cos^2x

OpenStudy (anonymous):

hold on:

OpenStudy (anonymous):

what happened to sin^2(x) in the last part???

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

you went from Sin^2(x) -1=Cos^2(x) to Cos^2(x)/Cos^2(x)=-1... idk how you got there

OpenStudy (anonymous):

sin^2(x) + cos^2(x) = 1 sin^2(x) + cos^2(x) - 1 = 0 -cos^2(x) -cos^2(x) sin^2(x) -1 = -cos^2(x)

OpenStudy (anonymous):

Do you know this trig identity?

OpenStudy (anonymous):

Are you still following? In this case, cos^2(x) is a common denominator, so you subtract the fraction like any other fraction. You subtract the numerator and keep the demominator of cos^2(x)